OpenStudy (babyslapmafro):

Please help me solve the following problem. Let z=3x^2-y^2. Find all points at which the magnitude of the gradient of z equals 6.

OpenStudy (babyslapmafro):

\[z=3x^2-y^2\]

terenzreignz (terenzreignz):

Let's take its gradient :D

OpenStudy (babyslapmafro):

<6x,-2y>

terenzreignz (terenzreignz):

oh... okay, great :D

terenzreignz (terenzreignz):

I suppose you also worked out its magnitude? :)

OpenStudy (babyslapmafro):

yea... do i find x and y in the following equation? \[x^2+\frac{ y^2 }{ 9 }=1\]

terenzreignz (terenzreignz):

OH, I see, were you TRYING to find specific values for x and y? :)

OpenStudy (babyslapmafro):

i looked in the back of the book and the answer is an equation so i guess I'm not supposed to find exact values

terenzreignz (terenzreignz):

No, in fact, I think this is enough, just add a little description :) From your vast precalculus experience, I hope you know that this equation represents an ellipse :D

OpenStudy (babyslapmafro):

The answer in the back of the book is : \[9x^2+y^2=9\] So my answer is essentially the same

OpenStudy (babyslapmafro):

yes i know it is an ellipse

terenzreignz (terenzreignz):

Well then, that was it :)

OpenStudy (babyslapmafro):

thanks

OpenStudy (babyslapmafro):

So a complete answer would be: The points at which the magnitude of the gradient of z equals 6 are contained on the ellipse given by the equation....(my original equation)

terenzreignz (terenzreignz):

Yes. (It IS an ellipse, right? Suddenly I'm not so sure of myself D: ) ^_^

OpenStudy (babyslapmafro):

ya

OpenStudy (babyslapmafro):